Determine the pressure drop along a venule of the length L and the radius R, if the viscosity of blood is ƞ and the speed of blood at the central layer of the venule is v0

1
Expert's answer
2021-04-06T18:25:24-0400

According to the Hagen–Poiseuille equation (see https://en.wikipedia.org/wiki/Hagen–Poiseuille_equation) the pressure drop is given as follows:


"\Delta P = \frac{8 \eta L Q}{\pi R^4}"

where "Q" is the volumetric flow rate. By definion, for a cylindrical pipe with only velocity component along the pipe, "Q" is given as follows:


"Q = v_0A"

where "A = \pi R^2" is the cross-sectional area of the pipe. Substituting the expression for "Q" in the first formula, obtain:


"\Delta P = \frac{8 \eta L v_0\pi R^2}{\pi R^4} = \frac{8 \eta L v_0}{R^2}"

Answer. "\Delta P = \frac{8 \eta L v_0}{R^2}"